OpenStudy (anonymous):

4(x + 1) = −12 last one, please help? :)

OpenStudy (anonymous):

4(x + 1) = -12 Distribute. 4x + 4 = -12 Subtract 4 from both sides. 4x = -16 Divide both sides by 4. x = -4

OpenStudy (anonymous):

I'm not sure how much you know the problem, so I'll talk about everything. Well first thing you need to do is distribute the -4 into the (x-1) We get -4x+4-12 Why I got this is because when you distribute you are multiplying the -4 by the "x" and the "-1" It becomes -4x+4 as shown above because -4 times "x" is -4x and -4 times -1 is 4 So we have -4x+4<12 We want to have "x" by itself on any side. The easiest side to have the "x" by itself is on the left side. So we want to minus 4 by both sides. The +4 will disappear and the 12 will become 8 Now its -4x+4<12 Now its -4x<8 Now we need to divide the equation by -4. Important! We also need to switch the sign because when you multiply or divide by a negative when dealing with this kind of equation.

OpenStudy (anonymous):

@ANAS23 4(x + 1) = −12 is an equation NOT an inequality.

OpenStudy (anonymous):

6 − x = 15 for this one the answer is 9 right? i tried doing it on my own but im not 100% positive its correct :/

OpenStudy (nubeer):

nop.. not 9

OpenStudy (anonymous):

6 − x = 15 Subtract 6. -x = 9 ^You have a -x which is the same as -1x. Divide both sides of the equation by -1.

OpenStudy (anonymous):

x = 9 /-1

OpenStudy (anonymous):

-9 ??

OpenStudy (anonymous):

Yes. (:

OpenStudy (anonymous):

seriously? 6 - x = 15 6 = 15 + x 6 - 15 = x -9 = x

OpenStudy (anonymous):

thank you :) i need help with one word problem but it sort of lengthy :/

OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

i understamd if you dont wanna help with that one lol

OpenStudy (anonymous):

I'll show you the steps. But you have to promise you'll explain it in your own words! 2(x + 6) − 10 = 12 + 4(x − 1) The FIRST thing you always check for are the parenthesis! Use the Distributive property. ~ \[2(x + 6) − 10 = 12 + 4(x − 1)\]\[2x + 12 - 10 = 12 + 4x - 4\] Combine like terms ~ *Remember to keep the sign in front of each term. \[2x + 12 - 10 = 12 + 4x - 4 \]\[2x + 2 = 8 + 4x \]Subtract 2x from both sides. \[2 = 8 + 2x \]Subtract 8 from both sides.\[-6 = 2x\]Divide both sides by 2.\[-3 = x\] So x = -3

OpenStudy (anonymous):

@Lyndsey_Coach_xox I hope this helps! (:

OpenStudy (anonymous):

Start by multiplying out the parentheses: 2(x + 6) − 10 = 12 + 4(x − 1) becomes: 2x + 12 - 10 = 12 + 4x - 4 Do the arithmetic to combine like terms to get: 2x + 2 = 4x + 8 Subtract 2x from both sides: 2 = 2x + 8 Subtract 8 from both sides: -6 = 2x Divide both sides by 2 and flip around: x = -3

OpenStudy (anonymous):

thank you soooooo much!!!

OpenStudy (anonymous):

i need my medal

OpenStudy (anonymous):

comment on my next question you dont have to answer it correct and ill give you one!

OpenStudy (anonymous):

Glad to help! (:

OpenStudy (anonymous):

im sorry but im finish the school in december